evaluate the function f(x)= 7x+7 at the given values if independent variable and simplify..
a. f(x+5)
b. f(-x)

thank you so much

Answer by ikleyn(52780)

.
evaluate the function f(x)= 7x+7 at the given values if independent variable and simplify..

a. f(x+5) = 7*(x+5) + 7 = (7x + 35) + 7 = 7x + (35+7) = 7x + 42.

b. f(-x) = 7*(-x) + 7 = -7x + 7.